Give a rough estimate (by rounding off to nearest hundreds) and also a closer estimate (by rounding off to nearest tens): 
489348 – 48365 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

489348–48365
Rounding off to nearest hundreds
489348–48365 = 489300–48400 = 440900
Rounding off to nearest tens
489348–48365 = 489350–48370 = 440980
​​
Thanks